What is the root of the equation x^2-x=6 (algebraically)?
Answer by ewatrrr(24785)


Hi,
Finding the roots
x^2-x-6 = 0
factoring
(x-3)(x+2)=0 Note:SUM of the inner product(-3x) and the outer product(2x) = -1x
(x-3)=0
x = 3
(x+2)=0
x = -2